# Non-universal surface magnetoelectric response in antiferromagnetic topological insulators

### v2

@@ Abstract
In Eq. (6) each 1D unit cell has four sites since there are two van der Waals layers, each with top and bottom surface Dirac cones, and $\lambda_{i}$ and $\tau_{i}$ are Pauli matrices that act on the layer and surface degrees of freedom. (Spin degree of freedom doesn’t appear in Eq. (6) since states in the $\mathcal{N}=0$ LL subspace are spin-up polarized.) The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (6) at half-filling: $P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{-\pi/2d}^{\pi/2d}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z})$, where $\xi^{a}_{nm}(k_{z})\equiv i\langle u_{n,k_{z}}|\frac{\partial}{\partial k^{z}}u_{m,k_{z}}\rangle$ is the Berry connection in a (periodic) Hamiltonian gauge. $P_{\text{SSH}}^{z}$ is quantized because Eq. (6) has particle-hole symmetry: $U^{\dagger}_{\mathcal{C}}(k_{z})H_{\text{AFM}}(-k_{z})^{*}U_{\mathcal{C}}(k_{z})=-H_{\text{AFM}}(k_{z})$ for $U_{\mathcal{C}}(k_{z})=(\lambda_{x}\text{cos}dk_{z}-\lambda_{y}\text{sin}dk_{z})\otimes\tau_{z}$. Indeed, at half-filling this model can support topologically distinct ground states characterized by a $\mathbb{Z}_{2}$-valued invariant  [^5] See Sec. 2D of Ref. [6] and Ref. [58]., which is obtained by performing the 1D BZ integral in $P^{z}_{\text{SSH}}$. The association  [^6] See Sec. 2D of Ref. [6]. of $\mathbb{Z}_{2}$-even with $P_{\text{SSH}}^{z}=en$ and $\mathbb{Z}_{2}$-odd with $P_{\text{SSH}}^{z}=e(n+\frac{1}{2})$ for some $n\in\mathbb{Z}$ allows Fig. 2 (a) to be confirmed by a semi-analytic calculation (see Supplementary Material). A similar analysis applies in the non-magnetic case, for which the projected Hamiltonian in the $\mathcal{N}=0$ subspace is the usual SSH model. Thus $P^{z}$ does not distinguish between the microscopically distinct magnetic-field dependent charge redistribution (Fig. 2) in AFM and non-magnetic cases.
We can also evaluate the bulk magnetoelectric coefficient in Eq. (1) using a 3D tight-binding Hamiltonian [15] obtained by regularizing Eq. (S1) and employing the previously derived [6, 7] linear response expression. As described above, in band insulators with (generalized) TRS it has been argued that $\alpha^{il}=\delta^{il}\alpha_{\text{CS}}$, where $\alpha_{\text{CS}}$ is sensitive to the TRS-induced topological classification of the electronic ground state (encoded in a $\mathbb{Z}_{2}$ index). In general, $\alpha_{\text{CS}}$ is gauge-dependent and its explicit form employs a smooth global gauge of the vector bundle of occupied Bloch states over the 3D BZ torus. Thus, a smooth global Hamiltonian gauge $\mathfrak{u}$ identified pointwise by $\mathfrak{u}_{\bm{k}}=(\ket{u_{n\bm{k}}})_{n\in\{1,\ldots,N_{\text{occ}}\}}$ is viable (if it exists) and in that gauge
@@
Explicitly computing $\alpha_{\text{CS}}$ is challenging because a smooth global Hamiltonian gauge does not typically exist in $\mathbb{Z}_{2}$-odd insulators [43, 44], requiring a non-trivial Wannierization-like procedure to be employed. In our case, however, because the model has a generalized TRS and all energy bands appear in isolated degenerate pairs over the entire BZ, such a gauge does exist  [^8] In general, TRS implies that the (set of) Chern number(s) associated with each set of isolated bands vanishes, thus there exists a smooth global frame of the corresponding vector bundles. Due to the degeneracy in this special case, the components of each such frame are energy eigenvectors. (This same argument holds if there are no degeneracies between bands anywhere in BZ, but in the presence of a symmetry such as time-reversal or inversion, which ultimately gives rise to a $\mathbb{Z}_{2}$ index, this will never be the case.) Hence, a smooth global Hamiltonian gauge always exists in such circumstances.. In a related work [15] we explicitly evaluated Eq. (7) in the non-magnetic case. We can, at the cost of some additional complication, also find the analytic form of a smooth global Hamiltonian gauge in the AFM case. With this gauge in-hand, and employing the previously detailed [15] approximation that the basis of the tight-binding model (that results from lattice regularization) consists of atomic-like Wannier functions, we can semi-analytically evaluate Eq. (7). (See Supplementary Material.) This calculation again reproduces quantizated $\alpha_{\text{CS}}$ assignments in agreement with Fig. 2 (a).
@@
Magnetoelectric response in thin films— In Fig. 3 (a) we show a $J_{\text{S}}=1.5\Delta_{\text{S}}$ phase diagram constructed by tracking thin film gap closings and confirmed by evaluating the magnetoelectric response coefficient $\alpha_{\text{me}}$, expressed here in units of $e^{2}/2hc$. The thin film results perfectly match the bulk $J_{\text{S}}=1.5\Delta_{\text{S}}$ results in Fig. 2. Notably, the thin film phase diagram has an extra boundary surrounding the region with $\alpha_{\text{me}}=2$, which results from the appearance of (top and bottom) surface Hall conductivities (Chern wrappers [14]) in this region of our model’s phase diagram, that are not related to bulk topology. The Chern wrapper region’s extent depends on details of the model that do not influence bulk topology; it is absent for $J_{\text{S}}\leq\Delta_{\text{S}}$, as illustrated in 3 (b). Note that the opportunity to identify bulk topological phase transitions by tracking thin film gaps is present only in magnetic $\mathbb{Z}_{2}$ TIs, which do not have gapless edge modes.
@@
Figure 3: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$. The (quantized) magnetoelectric coefficient $\alpha_{\text{me}}$ is expressed here in units of $e^{2}/2hc$. The green region corresponds to $\mathbb{Z}_{2}$-odd and the red and grey regions to $\mathbb{Z}_{2}$-even bulk phases; the states in the red region have Chern wrappers as discussed in the main text, and therefore have $\alpha_{\text{me}}=2$. (b) Thin film phase diagram for $J_{\text{S}}=0.8\Delta_{\text{S}}$. (c) Schematic illustration of ${\cal N}=0$ (see main text) charge distributions (red curves) for $\alpha_{\text{me}}=0,\,1,\,2$ at $J_{\text{S}}=1.5\Delta_{\text{S}}$ and a small value of $\Delta_{\text{D}}$. In all cases there are both interior and surface contributions to $\alpha_{\text{me}}$ that are not quantized separately. The arrows indicate the orientation of the static magnetic moments in each van der Waals layer and the background color (white, blue, green) indicates the layer Chern number in the limit $\Delta_{\text{D}}\to 0$.
@@
Figs. 3 (a) and (b) can be understood qualitatively by examining the $\Delta_{\text{D}}=0$ limit of the ${\cal N}=0$ Hamiltonian. In this hypothetical limit there is no hybridization between van der Waals layers and each layer is described by independent $2\times 2$ Hamiltonians with a surface degree-of-freedom:
@@
where $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$ with $m_{t/b}$ the exchange coupling at top and bottom surface of each layer. On the outermost surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$, with the sign specified by the moment orientation in the same layer [34]. The eigenvalues of this Hamiltonian are $E=\bar{m}\pm(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$. The phase diagram is controlled by the relative sizes of $|\bar{m}|$ and $(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ on the outermost and interior layers. In the $\alpha_{me}=0$ phase, $|\bar{m}|<(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, in the $\alpha_{me}=1$ phase $|\bar{m}|>(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, whereas in the $\alpha_{me}=2$ phase $|\bar{m}|$ is larger in the outside layers and smaller in the interior layers. In the $\alpha_{me}=1$ case, the ${\cal N}=0$ eigenfunctions occupy alternate layers as illustrated in the middle panel of Fig. 3 (c). For $\delta=0$ all layers are identical and
@@
is quantized even at finite $N$. More generally, for any values of $\Delta_{\text{D}}$ and $\delta$, the magnetoelectric coefficient achieves quantization only in sufficiently thick films [34].
@@
Discussion— We have explored the magnetoelectric properties of layered $\mathbb{Z}_{2}$ topological insulators by evaluating the (electronic) polarization response of realistic thin film and bulk model Hamiltonians to a uniform dc magnetic field along the stacking axis. For an effective low-energy model we find that the polarization response arises entirely from the formation of $\mathcal{N}=0$ anomalous Landau levels described by 1D generalized SSH models. In the thin film case the total magnetoelectric response is quantized for sufficiently thick films, but it arises from non-zero and non-universal surface and interior contributions. (This behaviour is partially related to that of the alternating stacked Haldane model considered in Ref. [46].) We have reached these conclusions both by numerically examining the many-layer limit of the thin film calculations and by applying periodic boundary conditions in the stacking direction to the SSH models. Both approaches find that only the total polarization is quantized, in agreement with our 3D bulk calculations that relate the magnetoelectric response to the Chern-Simons 3-form [6, 7], which we evaluate explicitly for a lattice regularization of the low-energy model.
@@
It is interesting to relate our results to efforts [47, 48] to express the the orbital magnetization ${\mathbf{M}}$ as a sum of contributions having either atomic-like ($\overline{{\mathbf{M}}}$) or itinerant ($\widetilde{{\mathbf{M}}}$) character. This partitioning can be made unambiguously, given a model with isolated energy bands that support localized Wannier functions, and associates $\overline{{\mathbf{M}}}$ with interior currents and $\widetilde{{\mathbf{M}}}$ with surface currents. However, when the bands are not isolated the partitioning is gauge-dependent. In previous work [15] on the non-magnetic case we were able, thanks to the analytic nature of our simplified model calculations, to demonstrate that there exists a smooth global (Hamiltonian) gauge $\mathfrak{u}$ in which the topological magnetoelectric response is entirely itinerant, i.e. $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}=\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. We conjectured that this property is not specific to the simplified model, but that it is instead a bulk manifestation of the feature that in non-magnetic thin films magnetoelectric response can occur only near the surface, and as a bulk harbinger of the magnetic-dopant requirement for the topological magnetoelectric effect to manifest in non-magnetic TI films. In the Supplementary Material we explicitly demonstrate that in the AFM case there does not generically exist a smooth global gauge in which $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}$ equals $\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. This difference between AFM and non-magnetic cases, which is encoded in the geometry  [^9] Recall that each global trivialization of a vector bundle corresponds with a global smooth frame (i.e., gauge) thereof (see, e.g., Chapter 10 of Ref. [LeeBook]). And although, for a given band insulator, the nonexistence of certain smooth global gauges has been related [44] to the topology of the vector bundle of occupied Bloch states $\mathcal{V}\rightarrow\text{BZ}$ [10], we do not suspect this to be the case here; we anticipate that in an AFM TI there could exist some smooth global gauges $\mathfrak{u}$ for which there exists a finely tuned point in the parameter space where ${\mathbf{M}}^{(E)}_{\mathfrak{u}}$ equals $\alpha_{\text{CS}}^{\mathfrak{u}}{\mathbf{E}}$. Thus, whether such a gauge generically exists is taken to be a geometric property of $\mathcal{V}$. of the corresponding vector bundles of occupied Bloch states, can be understood as a bulk manifestation of magnetoelectric response occurring both in the interior and at the surface of AFM TIs. Thus, although $\alpha_{\text{CS}}$ is determined by the (generalized) TRS-induced $\mathbb{Z}_{2}$ topological classification of the electronic ground state, which does not distinguish between non-magnetic and AFM TIs, the nature of that response – whether it is entirely itinerant or not – does distinguish them.
@@
Our results provide a conditional answer to the thorny conceptual issue as to whether the topological magnetoelectric effect should be viewed as an interior or surface effect. We have previously emphasized that the magnetoelectric response of ordinary non-magnetic TIs is localized at their surfaces, and is realized only when the surfaces are gapped by local magnetic dopants. When the bulk is magnetic, surface dopants are no longer required and a response in the interior is allowed by symmetry. Our finding that the interior response does occur but is not (separately) quantized, shows that it is the total response that is related to bulk topology, and that the mechanism by which it is manifested generically involves both interior and surface contributions. Indeed, we expect that this periodic microscopic scale magnetic-field-induced charge density variation that appears within AFM TIs (and is absent in non-magnetic TIs) can be observed using high-sensitivity high-resolution X-ray imaging techniques [50, 51, 52, 53, 54, 55], and that these measurements might be aided by the fact that the density varies only in the magnetic field direction. The inability to assign a quantized response of a topological state wholly to interior or surface is analogous to the related issue of non-universal interior-edge current partition [16, 17, 18, 19] in the IQHE. In both cases, bulk topology constrains the total response but not its partitioning.

### v4

@@ III Bulk phase diagram and polarization properties
In Eq. (6) each 1D unit cell has four sites since there are two van der Waals layers, each with top and bottom surface Dirac cones, and $\lambda_{i}$ and $\tau_{i}$ are Pauli matrices that act on the layer and surface degrees of freedom. (Spin degree of freedom does not appear in Eq. (6) since states in the $\mathcal{N}=0$ LL subspace are spin-up polarized.) The ground state polarization in our low-energy 3D model has one bulk component, namely $P^{z}$, which can be calculated using expressions derived in the modern theory of polarization [38, 39]. The $\mathcal{N}=0$ contribution to $P^{z}$ is proportional to the 1D bulk polarization of Eq. (6) at half-filling: $P_{\text{SSH}}^{z}=e\sum_{n=1}^{2}\int_{-\pi/2d}^{\pi/2d}\frac{dk_{z}}{2\pi}\xi^{z}_{nn}(k_{z})$, where $\xi^{z}_{nm}(k_{z})\equiv i\langle u_{n,k_{z}}|\frac{\partial}{\partial k^{z}}u_{m,k_{z}}\rangle$ is the Berry connection in a (periodic) Hamiltonian gauge. $P_{\text{SSH}}^{z}$ is quantized because Eq. (6) has particle-hole symmetry: $U^{\dagger}_{\mathcal{C}}(k_{z})H_{\text{AFM}}(-k_{z})^{*}U_{\mathcal{C}}(k_{z})=-H_{\text{AFM}}(k_{z})$ for $U_{\mathcal{C}}(k_{z})=(\lambda_{x}\text{cos}dk_{z}-\lambda_{y}\text{sin}dk_{z})\otimes\tau_{z}$. Indeed, at half-filling this model can support topologically distinct ground states characterized by a $\mathbb{Z}_{2}$-valued invariant  [^6] See Sec. 2D of Ref. [6] and Ref. [67]., which is obtained by performing the 1D BZ integral in $P^{z}_{\text{SSH}}$. The association  [^7] See Sec. 2D of Ref. [6]. of $\mathbb{Z}_{2}$-even with $P_{\text{SSH}}^{z}=en$ and $\mathbb{Z}_{2}$-odd with $P_{\text{SSH}}^{z}=e(n+\frac{1}{2})$ for some $n\in\mathbb{Z}$ allows Fig. 2(a) to be confirmed by a semianalytic calculation (see Appendix A.2). A similar analysis applies in the nonmagnetic case, for which the projected Hamiltonian in the $\mathcal{N}=0$ subspace is the usual SSH model. Thus, $P^{z}$ does not distinguish between the microscopically distinct magnetic field dependent charge redistribution (Fig. 2) in AFM and nonmagnetic cases.
We can also evaluate the bulk magnetoelectric coefficient in Eq. (1) using a 3D tight-binding Hamiltonian [15] obtained by regularizing Eq. (II) and employing the previously derived [6, 7] linear response expression. As described above, in band insulators with (generalized) TRS it has been argued that $\alpha^{il}=\delta^{il}\alpha_{\text{CS}}$, where $\alpha_{\text{CS}}$ is sensitive to the TRS-induced topological classification of the electronic ground state (encoded in a $\mathbb{Z}_{2}$ index). In general, $\alpha_{\text{CS}}$ is gauge dependent and its explicit form employs a smooth global gauge of the vector bundle of occupied Bloch states over the 3D BZ torus. Thus, a smooth global Hamiltonian gauge $\mathfrak{u}$ identified pointwise by $\mathfrak{u}_{\bm{k}}=(\ket{u_{n\bm{k}}})_{n\in\{1,\ldots,N_{\text{occ}}\}}$ is viable (if it exists) and in that gauge
@@
Explicitly computing $\alpha_{\text{CS}}$ is challenging because a smooth global Hamiltonian gauge does not typically exist in $\mathbb{Z}_{2}$-odd insulators [43, 44], requiring a nontrivial Wannierization-like procedure to be employed. In our case, however, because the model has a generalized TRS and all energy bands appear in isolated degenerate pairs over the entire BZ, such a gauge does exist  [^9] In general, TRS implies that the (set of) Chern number(s) associated with each set of isolated bands vanishes, thus there exists a smooth global frame of the corresponding vector bundles. Due to the degeneracy in this special case, the components of each such frame are energy eigenvectors. (This same argument holds if there are no degeneracies between bands anywhere in BZ, but in the presence of a symmetry such as time-reversal or inversion, which ultimately gives rise to a $\mathbb{Z}_{2}$ index, this will never be the case.) Hence, a smooth global Hamiltonian gauge always exists in such circumstances.. In a related work [15] we explicitly evaluated Eq. (7) in the nonmagnetic case. We can, at the cost of some additional complication, also find the analytic form of a smooth global Hamiltonian gauge in the AFM case. With this gauge in-hand, and employing the previously detailed [15] approximation that the basis of the tight-binding model (that results from lattice regularization) consists of atomic-like Wannier functions, we can semianalytically evaluate Eq. (7). (See Appendix B.1 and B.2.) This calculation again reproduces quantized $\alpha_{\text{CS}}$ assignments in agreement with Fig. 2(a).
IV Magnetoelectric response in thin films
@@
In Fig. 3(a) we show a $J_{\text{S}}=1.5\Delta_{\text{S}}$ phase diagram constructed by tracking thin film gap closings and confirmed by evaluating the magnetoelectric response coefficient $\alpha_{\text{me}}$, expressed here in units of $e^{2}/2hc$. The thin-film results perfectly match the bulk $J_{\text{S}}=1.5\Delta_{\text{S}}$ results in Fig. 2. Notably, the thin-film phase diagram has an extra boundary surrounding the region with $\alpha_{\text{me}}=2$, which results from the appearance of (top and bottom) surface Hall conductivities (Chern wrappers [14]) in this region of our model’s phase diagram, that are not related to bulk topology. The Chern wrapper region’s extent depends on details of the model that do not influence bulk topology; it is absent for $J_{\text{S}}\leq\Delta_{\text{S}}$, as illustrated in Fig. 3(b). Note that the opportunity to identify bulk topological phase transitions by tracking thin film gaps is present only in magnetic $\mathbb{Z}_{2}$ TIs, which do not have gapless edge modes.
@@
Figure 3: (a) Thin film phase diagram for $J_{\text{S}}=1.5\Delta_{\text{S}}$. The (quantized) magnetoelectric coefficient $\alpha_{\text{me}}$ is expressed here in units of $e^{2}/2hc$. The green region corresponds to $\mathbb{Z}_{2}$-odd and the red and gray regions to $\mathbb{Z}_{2}$-even bulk phases; the states in the red region have Chern wrappers as discussed in the main text, and therefore have $\alpha_{\text{me}}=2$. (b) Thin film phase diagram for $J_{\text{S}}=0.8\Delta_{\text{S}}$. (c) Schematic illustration of ${\cal N}=0$ (see main text) charge distributions (red curves) for $\alpha_{\text{me}}=0,\,1,\,2$ at $J_{\text{S}}=1.5\Delta_{\text{S}}$ and a small value of $\Delta_{\text{D}}$. In all cases there are both interior and surface contributions to $\alpha_{\text{me}}$ that are not quantized separately. The arrows indicate the orientation of the static magnetic moments in each van der Waals layer and the background color (white, blue, green) indicates the layer Chern number in the limit $\Delta_{\text{D}}\to 0$.
@@
Figures 3(a) and 3(b) can be understood qualitatively by examining the $\Delta_{\text{D}}=0$ limit of the ${\cal N}=0$ Hamiltonian. In this hypothetical limit there is no hybridization between van der Waals layers and each layer is described by independent $2\times 2$ Hamiltonians with a surface degree of freedom:
@@
where $\bar{m}\equiv(m_{t}+m_{b})/2$ and $\tilde{m}\equiv(m_{t}-m_{b})/2$ with $m_{t/b}$ the exchange coupling at the top and bottom surface of each layer. On the outermost surfaces $m_{t/b}=\pm J_{\text{S}}$ whereas on the interior surfaces $m_{t/b}=\pm(1-\delta)J_{\text{S}}$, with the sign specified by the moment orientation in the same layer (see Appendix A.3). The eigenvalues of this Hamiltonian are $E=\bar{m}\pm(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$. The phase diagram is controlled by the relative sizes of $|\bar{m}|$ and $(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ on the outermost and interior layers. In the $\alpha_{\text{me}}=0$ phase, $|\bar{m}|<(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, in the $\alpha_{\text{me}}=1$ phase $|\bar{m}|>(\Delta_{\text{S}}^{2}+\tilde{m}^{2})^{1/2}$ in all layers, whereas in the $\alpha_{\text{me}}=2$ phase $|\bar{m}|$ is larger in the outside layers and smaller in the interior layers. In the $\alpha_{\text{me}}=1$ case, the ${\cal N}=0$ eigenfunctions occupy alternate layers as illustrated in the middle panel of Fig. 3(c). For $\delta=0$ all layers are identical and
@@
is quantized even at finite $N$. More generally, for any values of $\Delta_{\text{D}}$ and $\delta$, the magnetoelectric coefficient achieves quantization only in sufficiently thick films (see Appendix A.3).
V Discussion
@@
We have explored the magnetoelectric properties of layered $\mathbb{Z}_{2}$ topological insulators by evaluating the (electronic) polarization response of realistic thin film and bulk model Hamiltonians to a uniform dc magnetic field along the stacking axis. For an effective low-energy model we find that the polarization response arises entirely from the formation of $\mathcal{N}=0$ anomalous Landau levels described by 1D generalized SSH models. In the thin film case the total magnetoelectric response is quantized for sufficiently thick films, but it arises from nonzero and nonuniversal surface and interior contributions. (This behavior is partially related to that of the alternating stacked Haldane model considered in Ref. [46].) We have reached these conclusions both by numerically examining the many-layer limit of the thin film calculations and by applying periodic boundary conditions in the stacking direction to the SSH models. Both approaches find that only the total polarization is quantized, in agreement with our 3D bulk calculations that relate the magnetoelectric response to the Chern-Simons 3-form [6, 7], which we evaluate explicitly for a lattice regularization of the low-energy model.
@@
It is interesting to relate our results to efforts [47, 48] to express the the orbital magnetization ${\mathbf{M}}$ as a sum of contributions having either atomic-like ($\overline{{\mathbf{M}}}$) or itinerant ($\widetilde{{\mathbf{M}}}$) character. This partitioning can be made unambiguously, given a model with isolated energy bands that support localized Wannier functions, and associates $\overline{{\mathbf{M}}}$ with interior currents and $\widetilde{{\mathbf{M}}}$ with surface currents. However, when the bands are not isolated the partitioning is gauge-dependent. In previous work [15] on the nonmagnetic case we were able, thanks to the analytic nature of our simplified model calculations, to demonstrate that there exists a smooth global (Hamiltonian) gauge $\mathfrak{u}$ in which the topological magnetoelectric response is entirely itinerant, i.e. $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}=\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. We conjectured that this property is not specific to the simplified model, but that it is instead a bulk manifestation of the feature that in nonmagnetic thin films magnetoelectric response can occur only near the surface, and as a bulk harbinger of the magnetic-dopant requirement for the topological magnetoelectric effect to manifest in nonmagnetic TI films. In Appendix B.3 we explicitly demonstrate that in the AFM case there does not generically exist a smooth global gauge in which $\widetilde{{\mathbf{M}}}_{\mathfrak{u}}^{(E)}$ equals $\alpha^{\mathfrak{u}}_{\text{CS}}{\mathbf{E}}$. This difference between AFM and nonmagnetic cases, which is encoded in the geometry  [^10] Recall that each global trivialization of a vector bundle corresponds with a global smooth frame (i.e., gauge) thereof (see, e.g., Chapter 10 of Ref. [68]). And although, for a given band insulator, the nonexistence of certain smooth global gauges has been related [44] to the topology of the vector bundle of occupied Bloch states $\mathcal{V}\rightarrow\text{BZ}$ [10], we do not suspect this to be the case here; we anticipate that in an AFM TI there could exist some smooth global gauges $\mathfrak{u}$ for which there exists a finely tuned point in the parameter space where ${\mathbf{M}}^{(E)}_{\mathfrak{u}}$ equals $\alpha_{\text{CS}}^{\mathfrak{u}}{\mathbf{E}}$. Thus, whether such a gauge generically exists is taken to be a geometric property of $\mathcal{V}$. of the corresponding vector bundles of occupied Bloch states, can be understood as a bulk manifestation of magnetoelectric response occurring both in the interior and at the surface of AFM TIs. Thus, although $\alpha_{\text{CS}}$ is determined by the (generalized) TRS-induced $\mathbb{Z}_{2}$ topological classification of the electronic ground state, which does not distinguish between nonmagnetic and AFM TIs, the nature of that response – whether it is entirely itinerant or not – does distinguish them.
@@
Our results provide a conditional answer to the thorny conceptual issue as to whether the topological magnetoelectric effect should be viewed as an interior or surface effect. We have previously emphasized that the magnetoelectric response of ordinary nonmagnetic TIs is localized at their surfaces, and is realized only when the surfaces are gapped by local magnetic dopants. When the bulk is magnetic, surface dopants are no longer required and a response in the interior is allowed by symmetry. Our finding that the interior response does occur but is not (separately) quantized, shows that it is the total response that is related to bulk topology, and that the mechanism by which it is manifested generically involves both interior and surface contributions. Indeed, we expect that this periodic microscopic scale magnetic-field-induced charge density variation that appears within AFM TIs (and is absent in nonmagnetic TIs) can be observed using high-sensitivity high-resolution x-ray imaging techniques [50, 51, 52, 53, 54, 55], and that these measurements might be aided by the fact that the density varies only in the magnetic field direction. The inability to assign a quantized response of a topological state wholly to interior or surface is analogous to the related issue of nonuniversal interior-edge current partition [17, 18, 19, 20] in the IQHE. In both cases, bulk topology constrains the total response but not its partitioning.
VI Acknowledgements